OpenStudy (anonymous):

cscxcotx(1-cos^2 x)=_____x

OpenStudy (campbell_st):

well start with csc = 1/sin cot = 1/tan = cos/sin 1 - cos^2 = sin^2 so its \[\frac{1}{\sin(x)} \times \frac{\cos(x)}{\sin(x)} \times \sin^2(x)\] just multiply it out and you'll have the answer